OpenStudy (anonymous):

I'm very lost, this problem has to do with solving linear equations. I need step by step please. A=L*W solve for L

OpenStudy (anonymous):

I know you have to switch it so it becomes L*W=A

Vocaloid (vocaloid):

"solve for W" means we want W by itself right now we have A = L*W, so we want to move L to the other side. we can do this by dividing both sides by L, giving us: A/L = W does that make sense?

OpenStudy (anonymous):

Not really :(

OpenStudy (anonymous):

so for solving L would i have to divide by w?

Vocaloid (vocaloid):

yes

Vocaloid (vocaloid):

if we want L by itself, we divide both sides by W A = L*W A/W = L

OpenStudy (anonymous):

oh okay